OpenStudy (anonymous):

Lydia received a parking ticket because she parked ten feet from a fire hydrant. She had estimated her distance from the fire hydrant as being the minimum allowable distance, which is fifteen feet. What was her percent error? fifteen percent thirty-three percent fifty percent sixty-seven percent

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Percent error = \(\frac {|Error|}{Theoritical Value} \times 100\)

OpenStudy (anonymous):

how long is this?

OpenStudy (anonymous):

\(|Error|= \) Experimental value \(-\) Theoretical value \(= |10-15| = 5 \) Substitute that in the equation and you get \(\frac {5}{15} \times 100 = 33\%\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome :)